OpenStudy (anonymous):

find the 1st and 2nd order of partial derivatives of f(x,y,z)= x^2yz^4e^z(y)^4

OpenStudy (across):

The format confuses me, is this what you mean? \[f(x,y,z)=x^2yz^4e^zy^4\]

OpenStudy (anonymous):

yes sorry, except at the end the z and y are to the power of e and y is to the power of 4

OpenStudy (across):

So, this: \[f(x,y,z)=x^2yz^4e^{zy^4}\]With respect to which variables, all three?

OpenStudy (anonymous):

yes all three

myininaya (myininaya):

treat everything else like a constant except the thing you are taking partial derivative with respect to

myininaya (myininaya):

i was looking at the first equation

myininaya (myininaya):

so we should be looking at the second equation right?

OpenStudy (anonymous):

yay partial derivatives are easy

myininaya (myininaya):

\[f_x=2x*yz^4e^{zy^4}\]

OpenStudy (anonymous):

yes that's the equation

OpenStudy (anonymous):

Sorry it's x^2 and everything else

myininaya (myininaya):

do you want to try f_z on your own

OpenStudy (anonymous):

cmon do it with respect to z

myininaya (myininaya):

\[f_y=x^2(1)z^4e^{zy^4}+x^2yz^4(4zy^3)e^{zy^4}\]

myininaya (myininaya):

those z's look funny to me

OpenStudy (across):

\[∂f/∂x=2xyz^4e^{zy^4}\]\[∂f/∂y=x^2z^4e^{zy^4}(4zy^4+1)\]\[∂f/∂z=x^2yz^3e^{zy^4}(zy^4+4)\]

OpenStudy (anonymous):

what about the second order of them?

OpenStudy (anonymous):

Does someone know how to do the second order?

myininaya (myininaya):

just do the same way we have been doing

myininaya (myininaya):

if you are taking derivative with respect to x then treat everything else like a constant except x

OpenStudy (anonymous):

ok, thank you all